OpenStudy (anonymous):

There are 2 points of questions here about linear eq..

OpenStudy (anonymous):

part a: Find a system of linear eq in the variables x1, x2, x3, whose solution set is given by the parametric eq x1=t; x2= 1+t; x3=2-t and I got the answer, please check... x1-x2 = -1 x3 + x1=2 part b: find another parametric solution to the system in part a in which the parameter is s and x3 =s I hv my answer but not sure can anyone help for part b?? thank youu

OpenStudy (ash2326):

@candyfloss94 you have three variables x1, x2 and x3 so you need to create three equation. You have made 2, find one more

OpenStudy (anonymous):

oh sorry, for part a it said,"find a system of TWO linear eq in the variables x1, x2, x3..." so I think my answer's enough, is it?

OpenStudy (ash2326):

oh then it's correct:)

OpenStudy (ash2326):

part b is also similar so x1=s x2=1+s ans x3=s

OpenStudy (anonymous):

so that's the parametric solutions you give??

OpenStudy (anonymous):

@ash2326 how about... because x3=s=2-t then x3=s x2=3-s x1=2-s which one is correct then? :/

OpenStudy (ash2326):

Sorry that was just the question for part b we have x1=s x2=1+s and x3=s

OpenStudy (ash2326):

Yeah you're right. I'm sorry, I got confused

OpenStudy (anonymous):

ah I see... that's alright. thanks so much for quick reply. I thought you've gone :o :p :)

OpenStudy (ash2326):

I was here:) Sorry and thanks:)